Prev Next

SW&IT Meeting Minutes 2022-03-28

Attendees: Brandon, David, Glenn, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20220321
LR20220321 - Bec, Matthieu 40.0
LR20220321 - Mechtley, Brandon 80.0
LR20220321 - Eychaner, Glenn 40.0
LR20220321 - Edmeades, David 40.0
LR20220321 - Grenz, Paul 0.0
LR20220321 - Hooper, Stephen 0.0
LR20220321 - Zhang, Yang 40.0
LR20220321 - Zhang, Xianyu 0.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

YZ

  • Previous:
    • Continued some work on MODS image checking program.
    • Other misc things on MODS/LBC.
    • Released a GCS patch for IT8315 (for John Hill's investigation).
    • TMS upcoming workshop discussion.
    • Archive update from Al (yes they probably won't be able to move on to our stuff for sometime, next meeting on Apr.10??).
    • Some work related the new OSS patch.
  • Next:
    • Need to prepare the talk for the upcoming TMS workshop.
    • Continue work on the MODS image checking program.
    • Continue some previous discussions with Olga on MODS/LBC issues (IT#8567 etc. and maybe more image checking/cleaning).
    • Maybe some work related on OSS patch.

DE

  • Last week:
    • Got a Veeam quote from CDW, in line with expectations ($2500 for 4 additional socket licenses and renewing our current maintenance agreement)
    • Heard from Ketiv, have deeper restore tests to do
    • Saw a couple more Veeam errors, will investigate with support to see what's going on.
    • Office trip to drop JW laptop off, verify SSD
    • Removed Vault Client 2018 from termserv; was confusing users, doesn't work anyway. Glad that they are able to use 2020 after getting sorted out.
    • AC computer arrived, he picked it up.
  • This week
    • Keep an eye on Jim; I don't think he's opened the new laptop yet
    • Re-restore Vault VM and test Vault actions on it while getting licensing from our prod VM
    • Veeam support work

GE

  • Last week:
    • Deployed LUCI 2.14.10, which fixes LMS file loading (7096) and MOS alignment (8604). Updated the RTD Problem Reporting document.
    • Helped D.Thompson look into 8603 (error setting detector config), but no conclusion was reached. The problem is not in the LUCI software, and may be a communication error between GEIRS and the ROE.
    • Created a luci_reset_aladin script which deletes the existing Aladin configuration and recreates it from scratch to resolve the problem where Aladin occasionally gets stuck in a weird state trying to access outside resources. Reported to Aladin developers.
    • Proposed the migration of the LUCI observer interface from Java/Swing to HTML/CSS/Javascript.
    • Found that moving LUCI from Java 8 (1.8) to Java 17 (latest long-term support), or in fact any more recent version of Java to try to improve performance is non-trivial due to changes in RMI (Activation) and XML (JAXB) support.
  • This week:
    • Continue bugfixes in the RTD.
    • Start looking at work on the TCS by talking with Yang and J.Hill.
    • Restart work on the observer GUI? Or start the migration to HTML/CSS/JS?

PSG

  • Last week
    • Fixed proxy issues with it resetting values to zero.
    • Fixed NOMIC initialization errors - camera now works, FITS header should be correct as well.
    • Added 'reconnect' to OSS so that if the proxy dies, the subsystem will re-establish communication.
    • Begun work on SHARK-NIR circular buffer configurability.
    • Re-did pull request to fix errors.
  • This week
    • Test OSS reconnect to proxy.
    • Test NOMIC with Steve Ertel.
    • Test SHARK-NIR code
    • Vacation starting on the 30th thru the 7th.

BM

  • Last week:
    • ao sw test planning meeting, updated test wiki pages (set_z0, software updates)
    • gain ramp updates from xianyu (more discussion, hold off on merge until tested)
    • interrupted offsets, discussion with juan carlos, investigation of logs (offload status)
    • more looking into msgd disconnects
  • This week
    • "investigations":
      • soul-sxwfs BCU timeouts
      • interrupted offsets
      • msgd disconnect
      • ncpa status rtdb: arcetri?
    • start playing with zeromq, dms re: ccdviewer

SH

  • Tucson network problems * continuing as of Saturday * different offices seem to be affected differently
  • Core switches
  • IT meeting * nothing much there
  • First run at 6-headed 507 computer

next:

  • Tucson PDU installation
  • New Synology - house LBTI data
  • AMD VMware install
Topic revision: r14 - 28 Mar 2022, MatthieuBec
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback